### v3

@@ 4.2 Model-based Indicators
where $P(x_{i})$ denotes the probability of $x_{i}$ occurring in the dataset. Later, Rényi entropy Rényi (1961) introduces an additional parameter $\alpha>0,\alpha\neq 1$ for a generalized entropy definition:
@@
Studies on biology and ecology Mouillot and Lepretre (1999); Peet (1974); He and Hu (2005); Gregorius and Gillet (2008) investigate Simpson’s Index (SI) Simpson (1949); Wu et al. (2024, 2022) for measuring the biodiversity of species and genetics. Zhou et al. (2020) proposes a variant of the original SI with a more flexible statistic metric:
@@
Vendi Score (VS) Dan Friedman and Dieng (2023); Pasarkar and Dieng (2023); Nguyen and Dieng (2024) is rencently proposed for diversity measurement in machine learning researches. Inspired by the Rényi entropy, a generalized VS metric Pasarkar and Dieng (2023) is defined as below:
@@
where $\bar{\lambda}_{i|\theta}$ denotes the normalized eigenvalues of the similarity kernel matrix $K_{S|\theta}$, and $supp(\bar{\lambda})$ is the set of indices of all non-zero eigenvalues. The smaller $\alpha<1$ makes the scoring more sensitive to rare classes and therefore allows accurate diversity measurement even under severe class imbalance. One simple implementation of the similarity kernel $K_{S|\theta}$ is to use the Gaussian Radial Basis function $k$ with feature embeddings as $k(g(x_{i}|\theta),g(x_{j}|\theta))=\exp(-\frac{1}{2}\|g(x_{i}|\theta)-g(x_{j}|\theta)\|^{2})$. Nguyen and Dieng (2024) further introduces quality scoring into Eq. 35 where for each subset $S_{b}\subset S$, its average quality score $Q(S_{b})=\frac{1}{|S_{b}|}\sum_{x_{i}\in S_{b}}IND_{i}$ is multiplied with $D^{VS}_{\alpha}(S_{b})$ for comprehensive evaluation in terms of quality and diversity.
Miranda et al. (2022) proposes an intrinsic diversity coefficient to measure the diversity of a dataset with Task2Vec embeddings Achille et al. (2019); Nguyen et al. (2019) for distance computation between different tasks. The Task2Vec encodes data from different tasks by the diagonal entries of the Fisher Information Matrix (FIM). The FIM results from fine-tuning only the final (e.g., token classification) layer of a pre-trained model, namely a probe model (e.g., GPT2 Radford et al. (2019)), to solve the task. Given a batch of samples $B$, the mathematical representation of FIM is defined as:
@@
where $\hat{x}_{i(j)}$ denotes the $j$-th token predicted from the model parameterized as $\theta$ given the real sequence input $x_{i(<j)}$. The expectation $\mathbbm{E}_{x_{i},j,\hat{x}_{i(j)}}$ takes an average over the sequence length $|x_{i}|$ for each $x_{i}$ sampled randomly from the batch $x_{i}\in B$. The Task2Vec embedding $\mathop{f_{B}}\limits^{\rightarrow}=diag(F_{B})$, where $diag(\cdot)$ denotes the diagonal entries of $F_{B}$. Based on the Task2Vec embeddings, Lee et al. (2023) proposes to compute the diversity coefficients $\hat{d}iv$ specifically for NLP datasets:
@@
Lu et al. (2023a) develops a diversity measure by open-ended tagging. Specifically, a tagging model parameterized by $\theta$ is trained with GPT4-labeled tagging pairs to describe each instruction tuning datapoint $x_{i}$ by its fine-grained, atomic intentions and semantics (e.g., tasks and domains). Correspondingly, the number of tags can be viewed as a diversity indicator for sampling a instruction subset $S_{b}$ from the whole set $S$ (see Alg. 1).
@@ 4.3 Geometry-based Coreset Sampling
Among different categories of coreset sampling methods, geometry-based methods are the most intuitive and widely-used ones Chen et al. (2012); Agarwal et al. (2020); Sener and Savarese (2017); Sinha et al. (2020); Kamalov (2020); Rezazadegan Tavakoli et al. (2011); Kirchenbauer et al. (2024); Zhou et al. (2023). The intuition behind is that close samples in the embedding space often share similar properties with low diversity. Therefore, redundant information can be effectively suppressed by controlling the minimum distance between any two samples for subset selection. Specifically, k-center greedy is a typical diversity-oriented sampling method for massive pretraining and instruction-tuning corpus Chen et al. (2023a); Bhatt et al. (2024); Wu et al. (2023); Zhao and Fang (2024); Du et al. (2023). It solves the minimax facility location (FL) problem Cornuéjols et al. (1983); Farahani and Hekmatfar (2009), i.e., selecting the subset $S_{b}$ under the given size budget $b$ from the full set $S$ so that the largest distance between an example in $S\backslash S_{b}$ and its closest example in $S_{b}$ is minimized:
@@
The direct solution to Eq. 38 is NP-hard Cook et al. (1994) and a greedy approximation is proposed Sener and Savarese (2017) (see Alg. 2).
Algorithm 2 K-Center Greedy Sener and Savarese (2017)
@@
For initialization of $S_{b}^{0}$, one can either choose randomly sampled datapoints from $S$, or use the cluster center points from $K$ clusters ($C_{1},C_{2},...,C_{K}$) of $S$ via k-means clustering. Similarly, the farthest point sampling method Eldar et al. (1997) shares the same principle that each iteration time only the farthest datapoint relative to the already selected coreset is chosen from the candidates.
In addition to the k-center greedy, the herding method Chen et al. (2012); Welling (2009); Huszár and Duvenaud (2012); Adhikary and Boots (2022) selects datapoints $x_{i}$ so that the distance between the coreset center and the full set center is minimized in the embedding space. For efficiency, it is also approximated via greedy implementation Chen et al. (2016); Harvey and Samadi (2014) by adding one sample each time into the $S_{b}$ to minimize the distance between two centers (see Alg. 3).
Algorithm 3 Herding Greedy Harvey and Samadi (2014)
@@
Furthermore, recent studies tend to develop complex heuristic sampling methods that takes geometry-based diversity into consideration Jiang et al. (2023c); Chan et al. (2021); Xia et al. (2022). Specifically, the inter-sample similarity of the selected coreset is minimized in return for an overall high diversity. Jiang et al. (2024c) proposes to preserve informative subset with the learning complexity (see Eq. 8) and implicitly puts constraints on its diversity via sampling on the k-means clusters:
@@
where $C$ denotes the constant that controls the degree of diversity. A larger $C$ represents the larger diversity of the dataset. The detailed procedure can be found in Alg. 4.
@@
Bukharin and Zhao (2023) proposes the quality-diversity instruction tuning (QDIT). It also uses FL functions for diversity measure of the subset $S_{b}$:
@@
where $sim(\cdot,\cdot)$ denotes the similarity function (e.g., cosine similarity). If the selected $S_{b}$ can be well-representative of the entire set $S$, then $S_{b}$ is assumed of high diversity. Given quality scores defined by Eq. 18, the detailed mechanism of QDIT is described in Alg. 5 with greedy approximation.
Algorithm 5 QDIT sampling Bukharin and Zhao (2023)
@@
Liu et al. (2023b) adopts the quality score-first and diversity-aware data selection method (DEITA), where all datapoints are first scored and sorted by quality measurement, and then selected by a geometry-based heuristic criterion (i.e., Repr Filter). Specifically, it considers that for each chosen datapoint in $S_{b}$, its $kNN_{i}^{1}$ (Eq. 26) should be above a certain threshold $\tau$ so that the overall diversity $D^{kNN}(S_{b})$ (Eq. 28) can be improved. As shown in Alg. 6, the quality and complexity of each sample $x_{i}$ is respectively measured by the trained complexity scoring model $\theta_{C}$ and the quality scoring model $\theta_{Q}$ with prompts $p_{C}$ and $p_{Q}$. Then, samples with high $G_{CQ}$ are prioritized but only those dissimilar ones can be kept for the diversity of $S_{b}$.
@@
Another series of geometry-based methods focus on the organization of data structures via developing clustering-based sampling techniques Citovsky et al. (2021); Tirumala et al. (2024); Axiotis et al. (2024); Shao et al. (2024); Alcoforado et al. (2024); Saranathan et al. . With respect to the clustering criterion, traditional methods employ topic modeling with LDA Blei et al. (2003); Raghuveer et al. (2012); Bui et al. (2017), NMF Lee and Seung (2000); Wang and Zhang (2012); Shen and Si (2010); Lazar and Doncescu (2009), TF-IDF Sparck Jones (1972); Bafna et al. (2016); Patil and Atique (2013); Roul et al. (2014), and latent concepts Xie et al. (2021) to assign text corpus into thematic clusters. Most recent studies exploit sentence encoding methods Reimers and Gurevych (2019); Feng et al. (2020) to perform clustering in the embedding space, where the vanilla k-means clustering and its variants Sinaga and Yang (2020); Kanungo et al. (2000); Bandyapadhyay and Varadarajan (2015), DBSCAN Deng (2020); Khan et al. (2014); Creţulescu et al. (2019), and spectral clustering Bach and Jordan (2003); Von Luxburg (2007); Jia et al. (2014) are widely used. Specifically, Tirumala et al. (2024) proposes to use SemDeDup Abbas et al. (2023) to remove semantically similar examples for deduplication, which provides a basis of diversity sampling. Then, k-means clustering is performed in the embedding space and prototype-based sampling technique Sorscher et al. (2022) is used. The “prototypical" samples, whose distance to their assigned cluster centers are small, should be discarded first to allow more “outliers" to be kept in $S_{b}$ during iterative sampling (see Alg. 7).
@@
Axiotis et al. (2024) proposed a k-means cluster-based sensitivity sampling technique. For each sample in one cluster, its distance to the cluster center and the proxy evaluation loss Feldman and Langberg (2011) of the center datapoint are both proportional to the probability of being chosed. Shao et al. (2024) proposes the balanced ClusterClip sampling. It first performs k-means clustering and then sample datapoints uniformly from each cluster. Different from the uniform sampling, the proposed ClusterClip puts constraints on the maximum number of each cluster being sampled, and therefore avoids overfitting of small clusters.
@@
Geometry-based sampling is intuitive and effective in diversity control. Most solutions to optimizing the overall diversity can be reformulated as variants of an iterative similarity or distance-based greedy sampling technique. Clustering does play an explanatory role in deciphering the embedding structures, making it easier and preciser to control the proportion of selection.
@@ 4.4 Bilevel Optimization-based Coreset Sampling
The selection of coreset can also be viewede as a bilevel optimization problem Colson et al. (2007); Zhang (2024); Sinha et al. (2017); Borsos et al. (2020); Killamsetty et al. (2021b, c); Zhang et al. (2022); Borsos et al. (2024); Pan et al. (2024) that consists of two loops: 1) the outer loop of optimizing the hard masks or soft weights for selecting the subset $S_{b}$ from $S$; 2) the inner loop of optimizing the model parameters $\theta$ on $S_{b}$. Without lose of generalizability, the bilevel optimization with the self-supervised language modeling loss can be written as follows:
$S_{b}^{*}$ $=\operatorname*{arg}\operatorname*{min}_{S_{b}\subset S}\sum_{x_{i}\in S_{b},\theta=\theta^{*}}NLL_{i}^{A|Q},$ (41)
@@
The retrieve method proposed by Killamsetty et al. (2021c) takes both labeled and unlabeled datasets into consideration, where the self-supervised loss from the unlabeled set (e.g., consistency regularization Xie et al. (2020); Wang et al. (2021c) and entropy regularization Zhao et al. (2020b); Grandvalet and Bengio (2004); Erkan and Altun (2010)) contributes to the inter-level and outer-level optimization as well. To improve the robustness, Glister Killamsetty et al. (2021b) optimizes the outer-level coreset selection on the additionally prepared validation set for the minimized validation loss. Li et al. (2023d) further emphasizes the role of the validation set in bilevel optimization. It not only computes the loss on the validation set for adversarial training, but also introduces gradient matching Killamsetty et al. (2021a) where the gradient of the model on the selected subset $S_{b}$ should be close to that on the entire $S$.
Borsos et al. (2024) reformulates the coreset sampling as a cardinality-constrained bilevel optimization problem. It proposes greedy forward selection and first-order methods that apply to any twice differentiable models. Variants of the solution for acceleration are extended: 1) binary weights, inverse-hessian-vector product approximations, and batch-wise selection; 2) small proxy models for fast estimation; 3) enforced sparsity-inducing penalty in the outer loop.
The ScaleBiO Pan et al. (2024) specifically addresses the data reweighting problem for large-scale LLM instruction tuning. It also prepares an extra validation set $S^{val}$ for the minimization of the outer loop. ScaleBio transforms the bilevel optimization into the single loop framework with an outer-level problem plus a constraint of the inner-level problem. A multiplier $\alpha>0$ and a proxy $u$ for optimizing the original inner loop (i.e., model weights $\theta$) are introduced into the minimax formulation Kwon et al. (2023); Lu and Mei (2024).
In contrast to a fixed budget $b$, Xia et al. (2024b) proposes a lexicographic bilevel-optimization method Borsos et al. (2020); Killamsetty et al. (2021b, c) where the inner loop optimizes model parameters and the outer loop optimizes data selection. When optimizing the selection mask, the minimization of loss terms is relaxed to allow the size of the final coreset smaller than $b$.
@@
The bilevel optimization methods often involve optimization regularization tricks as a relaxation to the original problem with nested outer-inner loops. Compared with the hard masks, the soft weights-based objective guarantees a higher level of diversity as each sample contributes more or less to the overall optimization.

### v4

@@ 4.2 Model-based Indicators
where $P(x_{i})$ denotes the probability of $x_{i}$ occurring in the dataset. Later, Rényi entropy Rényi (1961) introduce an additional parameter $\alpha>0,\alpha\neq 1$ for a generalized entropy definition:
@@
Studies on biology and ecology Mouillot & Lepretre (1999); Peet (1974); He & Hu (2005); Gregorius & Gillet (2008) investigate Simpson’s Index (SI) Simpson (1949); Wu et al. (2024; 2022) for measuring the biodiversity of species and genetics. Zhou et al. (2020) propose a variant of the original SI with a more flexible statistic metric:
@@
Vendi Score (VS) Dan Friedman & Dieng (2023); Pasarkar & Dieng (2023); Nguyen & Dieng (2024) is rencently proposed for diversity measurement in machine learning researches. Inspired by the Rényi entropy, a generalized VS metric Pasarkar & Dieng (2023) is defined as below:
@@
where $\bar{\lambda}_{i|\theta}$ denotes the normalized eigenvalues of the similarity kernel matrix $K_{S|\theta}$, and $supp(\bar{\lambda})$ is the set of indices of all non-zero eigenvalues. The smaller $\alpha<1$ makes the scoring more sensitive to rare classes and therefore allows accurate diversity measurement even under severe class imbalance. One simple implementation of the similarity kernel $K_{S|\theta}$ is to use the Gaussian Radial Basis function $k$ with feature embeddings as $k(g(x_{i}|\theta),g(x_{j}|\theta))=\exp(-\frac{1}{2}\|g(x_{i}|\theta)-g(x_{j}|\theta)\|^{2})$. Nguyen & Dieng (2024) further introduce quality scoring into Eq. 39 where for each subset $S_{b}\subset S$, its average quality score $Q(S_{b})=\frac{1}{|S_{b}|}\sum_{x_{i}\in S_{b}}IND_{i}$ is multiplied with $D^{VS}_{\alpha}(S_{b})$ for comprehensive evaluation in terms of quality and diversity.
Miranda et al. (2022) propose an intrinsic diversity coefficient to measure the diversity of a dataset with Task2Vec embeddings Achille et al. (2019); Nguyen et al. (2019) for distance computation between different tasks. The Task2Vec encodes data from different tasks by the diagonal entries of the Fisher Information Matrix (FIM). The FIM results from fine-tuning only the final (e.g., token classification) layer of a pre-trained model, namely a probe model (e.g., GPT2 Radford et al. (2019)), to solve the task. Given a batch of samples $B$, the mathematical representation of FIM is defined as:
@@
where $\hat{x}_{i(j)}$ denotes the $j$-th token predicted from the model parameterized as $\theta$ given the real sequence input $x_{i(<j)}$. The expectation $\mathbbm{E}_{x_{i},j,\hat{x}_{i(j)}}$ takes an average over the sequence length $|x_{i}|$ for each $x_{i}$ sampled randomly from the batch $x_{i}\in B$. The Task2Vec embedding $\mathop{f_{B}}\limits^{\rightarrow}=diag(F_{B})$, where $diag(\cdot)$ denotes the diagonal entries of $F_{B}$. Based on the Task2Vec embeddings, Lee et al. (2023) propose to compute the diversity coefficients $\hat{d}iv$ specifically for NLP datasets:
@@
Lu et al. (2023a) develop a diversity measure by open-ended tagging. Specifically, a tagging model parameterized by $\theta$ is trained with GPT4-labeled tagging pairs. To label such tags of each datapoint, the GPT4 first performs open-set fine-grained tagging and then all the collected tags are normalized to filter out low-frequency ones and aggregate near-duplicate ones. The specifically trained tagging model describes each instruction tuning datapoint $x_{i}$ by its fine-grained, atomic intentions and semantics (e.g., tasks and domains). Correspondingly, the number of tags can be viewed as a diversity indicator for sampling a instruction subset $S_{b}$ from the whole set $S$ (see Alg. 1).
@@ 4.3 Geometry-based Coreset Sampling
Among different categories of coreset sampling methods, geometry-based methods are the most intuitive and widely-used ones Chen et al. (2012); Agarwal et al. (2020); Sener & Savarese (2017); Sinha et al. (2020); Kamalov (2020); Rezazadegan Tavakoli et al. (2011); Kirchenbauer et al. (2024); Zhou et al. (2023). The intuition behind is that close samples in the embedding space often share similar properties with low diversity. Therefore, redundant information can be effectively suppressed by controlling the minimum distance between any two samples for subset selection. Specifically, k-center greedy is a typical diversity-oriented sampling method for massive pretraining and instruction-tuning corpus Chen et al. (2023a); Bhatt et al. (2024); Wu et al. (2023); Zhao & Fang (2024); Du et al. (2023). It solves the minimax facility location (FL) problem Cornuéjols et al. (1983); Farahani & Hekmatfar (2009), i.e., selecting the subset $S_{b}$ under the given size budget $b$ from the full set $S$ so that the largest distance between an example in $S\backslash S_{b}$ and its closest example in $S_{b}$ is minimized:
@@
The direct solution to Eq. 42 is NP-hard Cook et al. (1994) and a greedy approximation is proposed Sener & Savarese (2017) (see Alg. 2). Instead of simultaneously retrieving all the datapoints that can maximize the diversity of the selected subset, k-center greedy iteratively finds the most heterogeneous datapoint until the budget $b$ runs out. For initialization of $S_{b}^{0}$, one can either choose randomly sampled datapoints from $S$, or use the cluster center points from $K$ clusters ($C_{1},C_{2},...,C_{K}$) of $S$ via k-means clustering. Similarly, the farthest point sampling method Eldar et al. (1997) shares the same principle that each iteration time only the farthest datapoint relative to the already selected coreset is chosen from the candidates.
Algorithm 2 K-Center Greedy Sener & Savarese (2017)
@@
In addition to the k-center greedy, the herding methods Chen et al. (2012); Welling (2009); Huszár & Duvenaud (2012); Adhikary & Boots (2022) select datapoints $x_{i}$ so that the distance between the coreset center and the full set center is minimized in the embedding space. For efficiency, it is also approximated via greedy implementation Chen et al. (2016); Harvey & Samadi (2014) by adding one sample each time into the $S_{b}$ to minimize the distance between two centers (see Alg. 3). Similar to k-center greedy, the herding greedy also performs iterative sample selection. However, it differs in that in each step, the distance between the centers of the selected set and the full set is minimized.
Algorithm 3 Herding Greedy Harvey & Samadi (2014)
@@
Furthermore, recent studies tend to develop complex heuristic sampling methods that takes geometry-based diversity into consideration Jiang et al. (2023c); Chan et al. (2021); Xia et al. (2022). Specifically, the inter-sample similarity of the selected coreset is minimized in return for an overall high diversity. Jiang et al. (2024c) propose to preserve informative subset with the learning complexity (see Eq. 11) and implicitly puts constraints on its diversity via sampling on the k-means clusters:
@@
where $C$ denotes the constant that controls the degree of diversity. A larger $C$ represents the larger diversity of the dataset. The detailed procedure can be found in Alg. 4. It can be seen that datapoints are first sampled cluster-by-cluster to ensure a more balanced data distribution. Then, a percentile-based selection is performed on each cluster to select datapoints with easy learning complexity.
@@
Bukharin & Zhao (2023) propose the quality-diversity instruction tuning (QDIT). It also uses FL functions for diversity measure of the subset $S_{b}$:
@@
where $sim(\cdot,\cdot)$ denotes the similarity function (e.g., cosine similarity). If the selected $S_{b}$ can be well-representative of the entire set $S$, then $S_{b}$ is assumed of high diversity. Given quality scores defined by Eq. 21, the detailed mechanism of QDIT is described in Alg. 5 with greedy approximation. Different from the previous sequential setup that respectively prioritizes quality and diversity in two successive steps, QDIT adopts a dynamic weighting over the GPT scoring-based quality and the FL-based diversity for selection.
Algorithm 5 QDIT sampling Bukharin & Zhao (2023)
@@
Liu et al. (2023b) adopt the quality score-first and diversity-aware data selection method (DEITA), where all datapoints are first scored and sorted by quality measurement, and then selected by a geometry-based heuristic criterion (i.e., Repr Filter). Specifically, it considers that for each chosen datapoint in $S_{b}$, its $kNN_{i}^{1}$ (Eq. 30) should be above a certain threshold $\tau$ so that the overall diversity $D^{kNN}(S_{b})$ (Eq. 32) can be improved. As shown in Alg. 6, the quality and complexity of each sample $x_{i}$ are respectively measured by the trained complexity scoring model $\theta_{C}$ and the quality scoring model $\theta_{Q}$ with prompts $p_{C}$ and $p_{Q}$. Then, samples with high $G_{CQ}$ are prioritized but only those dissimilar ones can be kept for the diversity of $S_{b}$.
@@
Another series of geometry-based methods focus on the organization of data structures via developing clustering-based sampling techniques Citovsky et al. (2021); Tirumala et al. (2024); Axiotis et al. (2024); Shao et al. (2024); Alcoforado et al. (2024); Saranathan et al. . With respect to the clustering criterion, traditional methods employ topic modeling with LDA Blei et al. (2003); Raghuveer et al. (2012); Bui et al. (2017), NMF Lee & Seung (2000); Wang & Zhang (2012); Shen & Si (2010); Lazar & Doncescu (2009), TF-IDF Sparck Jones (1972); Bafna et al. (2016); Patil & Atique (2013); Roul et al. (2014), and latent concepts Xie et al. (2021) to assign text corpus into thematic clusters. Most recent studies exploit sentence encoding methods Reimers & Gurevych (2019); Feng et al. (2020) to perform clustering in the embedding space, where the vanilla k-means clustering and its variants Sinaga & Yang (2020); Kanungo et al. (2000); Bandyapadhyay & Varadarajan (2015), DBSCAN Deng (2020); Khan et al. (2014); Creţulescu et al. (2019), and spectral clustering Bach & Jordan (2003); Von Luxburg (2007); Jia et al. (2014) are widely used. Specifically, Tirumala et al. (2024) propose to use SemDeDup Abbas et al. (2023) to remove semantically similar examples for deduplication, which provides a basis of diversity sampling. Then, k-means clustering is performed in the embedding space. In each cluster, the prototype-based sampling technique Sorscher et al. (2022) is used. The "prototypical" samples, whose distance to their assigned cluster centers are small, should be discarded first to allow more "outliers" to be kept in $S_{b}$ during iterative sampling (see Alg. 7). It is noted that two times of clustering operations are conducted, where the semantic deduplication and prototype-based sampling are respectively performed on $K_{1}$ and $K_{2}$ clusters.
@@
Axiotis et al. (2024) propose a k-means cluster-based sensitivity sampling technique. For each datapoint in a cluster, both its distance to the cluster center and a proxy evaluation loss Feldman & Langberg (2011) measured on that cluster center contribute proportional to the probability of being chosen. Shao et al. (2024) propose the balanced ClusterClip sampling. It first performs k-means clustering and then sample datapoints uniformly from each cluster. Different from the uniform sampling, the proposed ClusterClip puts constraints on the maximum number of each cluster being sampled, and therefore avoids overfitting of small clusters.
@@
Geometry-based sampling is intuitive and effective in diversity control. Most solutions to optimizing the overall diversity can be reformulated as variants of an iterative, similarity or distance-based, greedy sampling technique. Clustering does play an explanatory role in deciphering the embedding structures, making it easier and preciser to control the proportion of selection.
@@ 4.4 Bilevel Optimization-based Coreset Sampling
The selection of coreset can also be viewede as a bilevel optimization problem Colson et al. (2007); Zhang (2024); Sinha et al. (2017); Borsos et al. (2020); Killamsetty et al. (2021b; c); Zhang et al. (2022); Borsos et al. (2024); Pan et al. (2024) that consists of two loops: 1) the outer loop of optimizing the hard masks or soft weights for selecting the subset $S_{b}$ from $S$; 2) the inner loop of optimizing the model parameters $\theta$ on $S_{b}$. Without lose of generalizability, the bilevel optimization with the self-supervised language modeling loss can be written as follows:
$S_{b}$ $=\operatorname*{arg}\operatorname*{min}_{S_{b}^{{}^{\prime}}\subset S}\sum_{x_{i}\in S_{b}^{{}^{\prime}},\theta=\theta^{*}}NLL_{i}^{A|Q},$ (45)
@@
The retrieve method proposed by Killamsetty et al. (2021c) takes both labeled and unlabeled datasets into consideration, where the self-supervised loss from the unlabeled set (e.g., consistency regularization Xie et al. (2020); Wang et al. (2021c) and entropy regularization Zhao et al. (2020b); Grandvalet & Bengio (2004); Erkan & Altun (2010)) contributes to the inter-level and outer-level optimization as well. To improve the robustness, Glister Killamsetty et al. (2021b) optimizes the outer-level coreset selection on the additionally prepared validation set for the minimized validation loss. Li et al. (2023e) further emphasize the role of the validation set in bilevel optimization. It not only computes the loss on the validation set for adversarial training, but also introduces gradient matching Killamsetty et al. (2021a) where the gradient of the model on the selected subset $S_{b}$ should be close to that on the entire $S$.
Borsos et al. (2024) reformulate the coreset sampling as a cardinality-constrained bilevel optimization problem. It proposes greedy forward selection and first-order methods that apply to any twice differentiable models. Variants of the solution for acceleration are extended: 1) binary weights, inverse-hessian-vector product approximations, and batch-wise selection; 2) small proxy models for fast estimation; 3) enforced sparsity-inducing penalty in the outer loop.
The ScaleBiO Pan et al. (2024) specifically address the data reweighting problem for large-scale LLM instruction tuning. It also prepares an extra validation set $S^{val}$ for the minimization of the outer loop. ScaleBio transforms the bilevel optimization into the single loop framework with an outer-level problem plus a constraint of the inner-level problem. A multiplier $\alpha>0$ and a proxy $u$ for optimizing the original inner loop (i.e., model weights $\theta$) are introduced into the minimax formulation Kwon et al. (2023); Lu & Mei (2024).
In contrast to a fixed budget $b$, Xia et al. (2024b) propose a lexicographic bilevel-optimization method Borsos et al. (2020); Killamsetty et al. (2021b; c) where the inner loop optimizes model parameters and the outer loop optimizes data selection. During optimization of the data selection mask, the loss terms are relaxed to allow the size of the final coreset smaller than $b$.
@@
The bilevel optimization methods often involve regularization tricks as a relaxation to the original problem with nested outer-inner loops. Compared with the boolean selection, soft weights-based objective guarantees a higher level of diversity as each sample contributes more or less to the overall optimization.